OpenStudy (anonymous):

can anyone do this? its a fraction ! 10x^4y^2 over -5x^3y^2 put the whole thing in parenthesis and an exponent of 3 outside

OpenStudy (anonymous):

\[\frac{10x^4y^2}{-5x^3y^2}=-2x^{4-3}=-2x\]

OpenStudy (anonymous):

and \[(-2x)^3=-8x^3\]